OpenStudy (anonymous):

24% converted to a fraction = ?

HanAkoSolo (jamierox4ev3r):

what is 24 percent a representative of? for example \(\frac{100}{100}\) is a representative of 100 %

OpenStudy (wolf1728):

24% = 24/100 =12/50 =6/25

OpenStudy (anonymous):

oh ok, I see.

HanAkoSolo (jamierox4ev3r):

yes it would basically be \(\frac{24}{100}\) or 24 out of 100... but that can be simplified down to \(\frac{12}{50}\) and then from there you can get to the final form of \(\frac{6}{25}\)

HanAkoSolo (jamierox4ev3r):

I kind of wanted @Zoobidoo to figure it out by themself @wolf1728 .... please make sure that happens next time :)

OpenStudy (anonymous):

Got it, thank you very much to you both. :)

OpenStudy (wolf1728):

As I've said here many times per cent literally means for every hundred.

HanAkoSolo (jamierox4ev3r):

right, a fraction is usually a number out of a hundred. Of course, it could also be out of a thousand if you are dealing with larger percentages... but I'm glad you understand :P

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Could I ask one more question please?

HanAkoSolo (jamierox4ev3r):

yes sure :)

OpenStudy (anonymous):

80% of 400 = ?

OpenStudy (anonymous):

Great, thank you.

OpenStudy (anonymous):

oh, so 80 x 4 = 320

HanAkoSolo (jamierox4ev3r):

yeah thats right :)

OpenStudy (anonymous):

Thank you. One more thing...

OpenStudy (anonymous):

for 60% of 390 would I do 390 x .60? or how else?

HanAkoSolo (jamierox4ev3r):

yes just search it you will be fine :) good luck!

OpenStudy (anonymous):

Thanks!